# Virtual Elastic Tether: a New Approach for Multi-agent Navigation in Confined Aquatic Environments

### v1

@@ IV Virtual Elastic Tether
where $\prescript{\mathcal{W}}{}{D}^{\mathcal{X}^{U}_{z=0}}_{\mathcal{X}^{S}}$ denotes the Euclidean distance between the projected trajectory of robot $U$ and robot $S$ on the plane $z=0$ at time $k$, while $\mathcal{F}$ denote the overall time-varying multi-objective controllers for the robots. Meanwhile, $\prescript{e}{}{\bm{u}}_{k}=\left[\prescript{e}{}{\bm{u}}^{U}_{k},\prescript{e}{}{\bm{u}}^{S}_{k}\right]^{\top}$ represent the control inputs for sub-tasks (such as depth control or tilt angles control) of robot $U$ and $S$, and $\prescript{\xi}{}{\bm{u}}_{k}=\left[\prescript{\xi}{}{\bm{u}}^{U}_{k},\prescript{\xi}{}{\bm{u}}^{S}_{k}\right]^{\top}$ represents the control inputs that drive the relative tether state $\xi_{k}=\left[\xi^{U\rightarrow S}_{k},\xi^{S\rightarrow U}_{k}\right]^{\top}$ towards its desired state $\xi_{d}=\left[\xi^{U\rightarrow S}_{d},\xi^{S\rightarrow U}_{d}\right]^{\top}$, which is user-specified value that can be set accordingly depending on a particular application. The constraint (IVc) accounts for the elasticity of the virtual tether.
@@
where $dim($·)))) denotes the dimensionality of the state vectors. The observation vector $\bm{z}^{U}=[z,\phi,\theta]^{\top}$ is supplemented with unobservable states $[x,y,\psi]^{\top}$ using $\bm{z}_{k}^{S}$ of robot $S$. Incorporating with (IV), the controller for the multi-agent system is given by
@@
Now, the remaining task is to design suitable functions $\mathcal{F}$, $E$, and $\Xi$ within the framework of (IVa-c). This feedback-based multi-agent formation framework can be adapted to various scenarios. For example, incorporating a local planner for obstacle avoidance into function $E$ transforms the system into a cooperative obstacle avoidance system with robust leader-follower formation capabilities. It is important to note that the measurements of $\xi$ are not restricted to any specific sensor type. They can be acoustic-based bearings, odometry-based distances, or visual-based relative positions, among others. The following section will provide an example of a visual-based VET implementation in an extreme aquatic environment where conventional systems fail to perform adequately.
@@ V Visual-based VET implementation of AUV-ASV Leader-follower Formation
In order to accomplish the underwater exploration mission in confined water ponds, where underwater communication and localisation are severely limited, this paper proposes a cooperative aquatic vehicle exploration system, abbreviated as CAVES, that incorporates both an AUV and an ASV. The system is comprised of two robots: the BlueROV2 AUV and the MallARD ASV [49]. While the AUV lacks self-localisation capabilities, the ASV can only localise itself on the water surface. As per (IVa,b), the BlueROV2 is equipped with a depth sensor and an IMU (Inertial Measurement Unit), while the MallARD utilises a 2-D LiDAR for surface localisation. The AUV is guided by the ASV, and their formation is stabilized by utilising VET to ensure robust guidance. Both robots are fitted with cameras, facing each other, to provide the measurement of the virtual tether state, $\xi$, as illustrated in Fig. 3.
@@ V-B Visual VET for Underwater Leader-Follower Formation
Following (IV) and (19), we define the error vector as
@@
Algorithm I:
LoS Awareness Mechanism
Input:
Position of the tag mounted on the robot in image frame $\prescript{\mathcal{C}}{}{\psi}$, $(\prescript{\mathcal{C}}{}{x}_{a},\prescript{\mathcal{C}}{}{y}_{a})$, $(\prescript{\mathcal{C}}{}{x}_{b},\prescript{\mathcal{C}}{}{y}_{b})$, $(\prescript{\mathcal{C}}{}{x}_{c},\prescript{\mathcal{C}}{}{y}_{c})$ and $(\prescript{\mathcal{C}}{}{x}_{d},\prescript{\mathcal{C}}{}{y}_{d})$
1:
If $\text{Detection}_{\text{tag}}=\text{True}$
2:
Calculate: $\bm{c}_{\text{tag}}\leftarrow(\prescript{\mathcal{C}}{}{\bar{x}},\prescript{\mathcal{C}}{}{\bar{y}})$, ${\bm{\dot{c}}}_{\text{tag}}\leftarrow(\frac{d\prescript{\mathcal{C}}{}{\bar{x}}}{dt},\frac{d\prescript{\mathcal{C}}{}{\bar{y}}}{dt})$
3:
If $\bm{c}_{\text{tag}}\in\Omega_{\text{safe}}$:
4:
  $[\prescript{\mathcal{C}}{}{u}_{k|x},\prescript{\mathcal{C}}{}{u}_{k|y}]^{\top}=K_{s|p}[\bm{c_{\text{tag}}}-\bm{c}_{\text{image}}]^{\top}$
5:
Else if $\bm{c}_{\text{tag}}\in\Omega_{\text{elastic}}$ :
6:
  $[\prescript{\mathcal{C}}{}{u}_{k|x},\prescript{\mathcal{C}}{}{u}_{k|y}]^{\top}=K_{e|p}[\bm{c_{\text{tag}}}-\bm{c}_{\text{image}}]^{\top}+K_{e|d}\bm{\dot{c}_{\text{tag}}}$
7:
Else:
8:
  $[\prescript{\mathcal{C}}{}{u}_{k|x},\prescript{\mathcal{C}}{}{u}_{k|y}]^{\top}=[\prescript{\mathcal{C}}{}{u}_{\text{max}|x},\prescript{\mathcal{C}}{}{u}_{\text{max}|y}]^{\top}$
9:
End if
@@
11:
$\prescript{\mathcal{C}}{}{u}_{k|\psi}=K_{\psi}\prescript{\mathcal{C}}{}{\psi}_{k}$
Output:
Control input $\prescript{\mathcal{C}}{}{u}_{k}$
In function $\Xi$, the proportional gains meet $K_{s|p}<K_{e|p}$, the max input $[\prescript{\mathcal{C}}{}{u}_{max|x},\prescript{\mathcal{C}}{}{u}_{max|y}]^{\top}$ is determined by the hardware constraints of the robot, it meets $[\prescript{\mathcal{C}}{}{u}^{U}_{max|x},\prescript{\mathcal{C}}{}{u}^{U}_{max|y}]^{\top}=[\prescript{\mathcal{C}}{}{u}^{S}_{max|x},\prescript{\mathcal{C}}{}{u}^{S}_{max|y}]^{\top}$, corresponding to assumption 1 (5). Therefore, when the desired tether state $\xi_{\text{des}}$ is achieved, such that $\bm{c_{\text{tag}}}=\bm{c}_{\text{image}}$, and recalling the feedback loop defined in (IV), we can then formulate the function $\Xi^{U}$ and $\Xi^{S}$ as follows:
@@ VI-A1 Objective & setup
Figure 7: In the results of the convergence experiment, the time-varying 6DoF velocities of both robots are displayed in the top two figures. The bottom figure shows the measured tether states (as per (IV), with $\xi_{d}=0$) relative to each robot. Both robots’ velocities and relative tether states converge to values close to zero, demonstrating that the proposed system is controllable in aquatic environments in a nominal scenario.

### v2

@@ IV Virtual Elastic Tether
where $\prescript{\mathcal{W}}{}{D}^{\mathcal{X}^{U}_{z=0}}_{\mathcal{X}^{S}}$ denotes the Euclidean distance between the projected trajectory of robot $U$ and robot $S$ on the plane $z=0$ at time $k$, while $\mathcal{F}$ denote the overall time-varying multi-objective controllers for the robots. Meanwhile, $\prescript{e}{}{\bm{u}}_{k}=\left[\prescript{e}{}{\bm{u}}^{U}_{k},\prescript{e}{}{\bm{u}}^{S}_{k}\right]^{\top}$ represent the control inputs for sub-tasks (such as depth control or tilt angles control) of robot $U$ and $S$, and $\prescript{\xi}{}{\bm{u}}_{k}=\left[\prescript{\xi}{}{\bm{u}}^{U}_{k},\prescript{\xi}{}{\bm{u}}^{S}_{k}\right]^{\top}$ represents the control inputs that drive the relative tether state $\xi_{k}=\left[\xi^{U\rightarrow S}_{k},\xi^{S\rightarrow U}_{k}\right]^{\top}$ towards its desired state $\xi_{d}=\left[\xi^{U\rightarrow S}_{d},\xi^{S\rightarrow U}_{d}\right]^{\top}$, which is user-specified value that can be set accordingly depending on a particular application. The constraint (10c) accounts for the elasticity of the virtual tether.
@@
where $dim($·)))) denotes the dimensionality of the state vectors. The observation vector $\bm{z}^{U}=[z,\phi,\theta]^{\top}$ is supplemented with unobservable states $[x,y,\psi]^{\top}$ using $\bm{z}_{k}^{S}$ of robot $S$. Incorporating with (11), the controller for the multi-agent system is given by
@@
Now, the remaining task is to design suitable functions $\mathcal{F}$, $E$, and $\Xi$ within the framework of (14a-c). This feedback-based multi-agent formation framework can be adapted to various scenarios. For example, incorporating a local planner for obstacle avoidance into function $E$ transforms the system into a cooperative obstacle avoidance system with robust leader-follower formation capabilities. It is important to note that the measurements of $\xi$ are not restricted to any specific sensor type. They can be acoustic-based bearings, odometry-based distances, or visual-based relative positions, among others. The following section will provide an example of a visual-based VET implementation in an extreme aquatic environment where conventional systems fail to perform adequately.
@@ V Visual-based VET implementation of AUV-ASV Leader-follower Formation
In order to accomplish the underwater exploration mission in confined water ponds, where underwater communication and localisation are severely limited, this paper proposes a cooperative aquatic vehicle exploration system, abbreviated as CAVES, that incorporates both an AUV and an ASV. The system is comprised of two robots: the BlueROV2 AUV and the MallARD ASV [49]. While the AUV lacks self-localisation capabilities, the ASV can only localise itself on the water surface. As per (9a,b), the BlueROV2 is equipped with a depth sensor and an IMU (Inertial Measurement Unit), while the MallARD utilises a 2-D LiDAR for surface localisation. The AUV is guided by the ASV, and their formation is stabilized by utilising VET to ensure robust guidance. Both robots are fitted with cameras, facing each other, to provide the measurement of the virtual tether state, $\xi$, as illustrated in Fig. 3.
@@ V-B Visual VET for Underwater Leader-Follower Formation
Following (8) and (19), we define the error vector as
@@
Algorithm I: LoS Awareness Mechanism
Input: Position of the tag mounted on the robot in image frame $\prescript{\mathcal{C}}{}{\psi}$, $(\prescript{\mathcal{C}}{}{x}_{a},\prescript{\mathcal{C}}{}{y}_{a})$, $(\prescript{\mathcal{C}}{}{x}_{b},\prescript{\mathcal{C}}{}{y}_{b})$, $(\prescript{\mathcal{C}}{}{x}_{c},\prescript{\mathcal{C}}{}{y}_{c})$ and $(\prescript{\mathcal{C}}{}{x}_{d},\prescript{\mathcal{C}}{}{y}_{d})$
1: If $\text{Detection}_{\text{tag}}=\text{True}$
2: Calculate: $\bm{c}_{\text{tag}}\leftarrow(\prescript{\mathcal{C}}{}{\bar{x}},\prescript{\mathcal{C}}{}{\bar{y}})$, ${\bm{\dot{c}}}_{\text{tag}}\leftarrow(\frac{d\prescript{\mathcal{C}}{}{\bar{x}}}{dt},\frac{d\prescript{\mathcal{C}}{}{\bar{y}}}{dt})$
3: If $\bm{c}_{\text{tag}}\in\Omega_{\text{safe}}$:
4:   $[\prescript{\mathcal{C}}{}{u}_{k|x},\prescript{\mathcal{C}}{}{u}_{k|y}]^{\top}=K_{s|p}[\bm{c_{\text{tag}}}-\bm{c}_{\text{image}}]^{\top}$
5: Else if $\bm{c}_{\text{tag}}\in\Omega_{\text{elastic}}$ :
6:   $[\prescript{\mathcal{C}}{}{u}_{k|x},\prescript{\mathcal{C}}{}{u}_{k|y}]^{\top}=K_{e|p}[\bm{c_{\text{tag}}}-\bm{c}_{\text{image}}]^{\top}+K_{e|d}\bm{\dot{c}_{\text{tag}}}$
7: Else:
8:   $[\prescript{\mathcal{C}}{}{u}_{k|x},\prescript{\mathcal{C}}{}{u}_{k|y}]^{\top}=[\prescript{\mathcal{C}}{}{u}_{\text{max}|x},\prescript{\mathcal{C}}{}{u}_{\text{max}|y}]^{\top}$
9: End if
@@
11: $\prescript{\mathcal{C}}{}{u}_{k|\psi}=K_{\psi}\prescript{\mathcal{C}}{}{\psi}_{k}$
Output: Control input $\prescript{\mathcal{C}}{}{u}_{k}$
In function $\Xi$, the proportional gains meet $K_{s|p}<K_{e|p}$, the max input $[\prescript{\mathcal{C}}{}{u}_{max|x},\prescript{\mathcal{C}}{}{u}_{max|y}]^{\top}$ is determined by the hardware constraints of the robot, it meets $[\prescript{\mathcal{C}}{}{u}^{U}_{max|x},\prescript{\mathcal{C}}{}{u}^{U}_{max|y}]^{\top}=[\prescript{\mathcal{C}}{}{u}^{S}_{max|x},\prescript{\mathcal{C}}{}{u}^{S}_{max|y}]^{\top}$, corresponding to assumption 1 (5). Therefore, when the desired tether state $\xi_{\text{des}}$ is achieved, such that $\bm{c_{\text{tag}}}=\bm{c}_{\text{image}}$, and recalling the feedback loop defined in (11), we can then formulate the function $\Xi^{U}$ and $\Xi^{S}$ as follows:
@@ VI-A1 Objective & setup
Figure 7: In the results of the convergence experiment, the time-varying 6DoF velocities of both robots are displayed in the top two figures. The bottom figure shows the measured tether states (as per (10), with $\xi_{d}=0$) relative to each robot. Both robots’ velocities and relative tether states converge to values close to zero, demonstrating that the proposed system is controllable in aquatic environments in a nominal scenario.